Find θ if θ is between 0° and 90°. Round your answer to the nearest tenth of a degree.

cos θ = 0.8890

Find θ if θ is between 0° and 90°. Round your answer to the nearest tenth of a degree.

sin θ = 0.9831

Find θ if θ is between 0° and 90°. Round your answer to the nearest tenth of a degree.

csc θ = 1.6195
Answer #1

Soluno · Givendo → - coso=0.88.90 A = cos (0.8890) o = 27.2 o Given. Sine = 0.9831 → 0 = sin (0.9831) + 0 = 79,45 · Giveni CS